MichaelWilliams4:

Gina looked at the architectural plan of a room with four walls in which the walls meet each other at right angles. The length of one wall in the plan is 17 inches. The length of the diagonal of the floor of the room in the plan is approximately 18.79 inches. Is the room in the shape of a square? Explain how you determined your answer. Show all your work.

AZ:

Hey there! Welcome to QuestionCove! Let's first draw the image out 

AZ:

For it to be a square, that means all FOUR sides have to be equal that means that 'x' has to be 17 do you know how to solve for x? We have a right triangle so we can use Pythagorean theorem a^2 + b^2 = c^2 where a and b are just two sides of the triangle and c is the hypotenuse (or the longest side- this is the side opposite the 90 degree angle) so if we set up the formula for your question x^2 + 17^2 = 18.79^2 can you solve for 'x'